# Mantis: A Simulation-Grounded Foundation Model for Disease Forecasting

### v3

@@ Abstract
Infectious disease forecasting in novel outbreaks or low-resource settings is hampered by the need for disease-specific data, bespoke training, and expert tuning. We introduce Mantis, a foundation model trained entirely on mechanistic simulations, which enables out-of-the-box forecasting across diseases, regions, and outcomes, even in settings with limited historical data. We evaluated Mantis against 48 forecasting models across six diseases with diverse transmission modes, assessing both point forecast accuracy (mean absolute error) and probabilistic performance (weighted interval score and coverage). Despite using no real-world data during training, Mantis achieved lower mean absolute error than all models in the CDC’s COVID-19 Forecast Hub when backtested on early pandemic forecasts. Across all other diseases tested, including respiratory, vector-borne, and waterborne pathogens, Mantis consistently ranked in the top two models across all evaluation metrics. Notably, Mantis generalized to diseases with transmission mechanisms not represented in its training data, demonstrating that it captures fundamental contagion dynamics rather than memorizing disease-specific patterns. These capabilities position Mantis as a practical foundation for disease forecasting: general-purpose, accurate, and deployable where traditional models fail.
@@ Introduction
Accurate forecasting of infectious disease dynamics is fundamental to effective public health response. During the COVID-19 pandemic, forecasts touched nearly every aspect of daily life—from how scarce hospital beds were allocated to if restaurants could open and whether families could send their children to school [1]. Yet despite their importance, disease forecasts often fail in practice, particularly when they are most needed. During the first two years of the COVID-19 pandemic, the performance of many models in the CDC Forecast Hub was worse than a naïve baseline that simply projected the most recent value forward [2]. This highlights a key limitation of many operational systems: they struggle to provide accurate, actionable forecasts precisely when uncertainty is highest and decisions are most urgent.
@@
Several structural limitations contribute to this breakdown. First, most models require large volumes of disease-specific historical data, which are rarely available early in outbreaks. Second, they rely on extensive manual tuning by expert teams, limiting scalability and access in under-resourced health departments, such as those in rural counties or low- and middle-income countries. Third, they generalize poorly across diseases, locations, or targets, necessitating new models for every context. In addition, their forecasting horizons are typically short: most models only predict 1–4 weeks ahead [2, 3, 4], far too short-term for real-world public health planning. Key decisions, such as ramping up hospital capacity, acquiring ventilators, or coordinating vaccination campaigns, often require 6–8 weeks of lead time. Without reliable long-range forecasts, these decisions are either delayed or made reactively, often at considerable cost [5].
@@
Typically, projections beyond 4 weeks are restricted to scenario analysis, which explores hypothetical outcomes conditional on specific policy interventions or behavioral changes [6, 7]. However, unconditional forecasts at longer horizons remain valuable for public health planning. We define forecasts as unconditional projections of expected outcomes, versus scenario analyses that specify hypothetical interventions. By this definition, temporal horizon does not determine whether a projection is a forecast or a scenario—an 8-week unconditional projection is a forecast, while a 2-week conditional projection is still scenario analysis. Public health decisions such as hospital capacity expansion and vaccination campaigns often require 6-8 weeks of lead time, motivating the development of models capable of reliable longer-range forecasting without any policy assumptions.
Addressing these challenges (limited historical data, poor cross-disease generalization, restricted forecast horizons, and the need for extensive expert tuning) requires a different approach to model training. Simulation-Grounded Neural Networks (SGNNs) provide such a framework [8]. SGNNs learn entirely from large-scale, highly realistic mechanistic simulations, using data generated by domain-specific simulators to train flexible neural networks. In epidemiology, this means generating synthetic outbreaks from mechanistic disease models that capture processes such as transmission, progression, immunity, seasonality, and intervention effects, then training models on these simulated trajectories so they can forecast real-world outbreaks they have never encountered before. By grounding learning in simulated systems that encode known mechanisms—whether physical, biological, or social—SGNNs enable models to generalize to real-world settings where data are limited, noisy, or missing.
@@
We present Mantis, the first large-scale foundation model built on the SGNN paradigm. Building on early demonstrations of simulation-grounded forecasting [9], Mantis is trained entirely on mechanistic simulations without access to any real-world training data. It learns from over 400 million simulated days of outbreak dynamics spanning diverse pathogens, transmission routes (human-to-human, vector-borne, environmental), intervention strategies, immunity profiles, and surveillance artifacts such as underreporting and delays. These simulations encode core epidemiological mechanisms, allowing Mantis to internalize generalizable contagion dynamics rather than extrapolating from historical trends. Unlike traditional models, Mantis requires no disease-specific tuning or retraining, yet it produces accurate probabilistic forecasts out-of-the-box across diseases, geographies, and targets (Figure 1).
@@
Figure 1: Conceptual overview of Mantis. Mantis is a simulation-grounded foundation model trained entirely on synthetic outbreaks generated by mechanistic epidemiological models. The training pipeline begins with a modular simulator that encodes diverse outbreak mechanisms, including multiple transmission modes (human-to-human, vectorborne, environmental), progression dynamics, intervention strategies, and population structures. These simulations are then passed through an observation model that incorporates real-world surveillance effects such as underreporting, stochasticity, and reporting delays. Trained on over 400 million simulated days, Mantis forecasts directly from real-world time series at inference time. Key advantages include accurate out-of-the-box performance, long-range forecasting capability, and mechanistic interpretability via back-to-simulation attribution.
@@
We evaluate Mantis against 48 models across forecasting tasks spanning six real-world diseases: seasonal influenza, early COVID-19, dengue, hepatitis B, smallpox, and scarlet fever. These evaluations cover multiple transmission modes, outbreak phases, surveillance systems, and historical eras. We assessed both point forecast accuracy and probabilistic performance. When benchmarked against CDC Forecast Hub models on early pandemic COVID-19 forecasts, Mantis achieved the lowest mean absolute error despite using no COVID-specific data or covariates in training and was tied for the second lowest weighted interval score. Across all other diseases—including those with transmission mechanisms absent from its training data—Mantis consistently ranked among the top two models on all evaluation metrics, demonstrating robust generalization to diverse epidemiological settings.
@@
These capabilities establish Mantis as a general-purpose forecasting system offering a scalable solution for modeling in data-scarce or rapidly evolving settings. To support deployment, we release Mantis as an open-source repository. Users can generate forecasts from a single historical time series without any machine learning expertise. 222github.com/carsondudley1/Mantis: Pretrained models, example data, and an interactive forecasting demo are available. Mantis is currently contributing forecasts to a new local-scale forecasting hub within Michigan for the 2025-2026 respiratory virus season and will be submitted to the CDC Forecast Hub to support real-time public health decision-making.
@@ Results
We compare Mantis against a range of published and re-implemented models, including top submissions to the CDC COVID-19 Forecast Hub, statistical baselines (exponential smoothing and SARIMA), LSTM neural networks retrained at each forecast point, and Chronos, a state-of-the-art time-series foundation model [18].
Mantis achieved competitive performance across diseases and metrics (Table 1). For early pandemic COVID-19 forecasts at the state level, Mantis achieved the lowest MAE among evaluated models, including the CDC Forecast Hub ensemble that combines dozens of expert-tuned submissions. The CDC ensemble achieved slightly better WIS (0.61 vs 0.63), while UMass-MechBayes achieved marginally better 95% coverage (0.93 vs 0.91), demonstrating that multiple approaches can excel on different aspects of probabilistic forecasting. Although results have shown that some general time series forecasting foundation models can confer reasonable results for infectious disease forecasting [19], we did not find that any we tested gave competitive performance with Mantis or our other baseline models, so we did not include them in our analysis.
@@
Example forecast trajectories illustrate Mantis’s ability to match the timing and magnitude of real-world outbreaks across diverse epidemiological settings (Figure 3). Forecasts include uncertainty intervals derived from the model’s predicted quantiles, enabling direct use in probabilistic planning. Because Mantis is trained entirely on broad mechanistic simulations rather than historical observations, it requires no model fitting or fine-tuning for new targets.
@@
State-level COVID Mantis 0.65 0.63 0.51 0.91
@@
Local-scale flu hosp. Mantis 0.85 0.86 0.51 0.92
@@
Dengue cases Mantis 0.75 0.80 0.47 0.90
@@
Smallpox cases Mantis 0.83 0.84 0.45 0.87
@@
Scarlet Fever cases Mantis 0.80 0.81 0.47 0.91
@@
Hepatitis B cases Mantis 0.85 0.84 0.47 0.90
@@
We evaluated two variants of Mantis to assess the impact of additional information on forecasting performance: a covariate-free version, which receives only the time series to be forecast (e.g., deaths), and a covariate-aware version, which also receives related signals (e.g., cases or hospitalizations). This design mirrors real-world forecasting settings, where models may have access to upstream indicators with predictive value. (See Methods for training details for each model.)
For example, when forecasting COVID-19 deaths, the covariate-aware Mantis version receives both historical deaths and upstream case data, while the covariate-free version sees only the death time series. Both models use identical architecture and training data; the only difference is the presence or absence of additional input channels.
@@
Figure 2: Covariate integration improves accuracy. Mantis maintains calibrated uncertainty across forecast horizons. (a) Including covariates (e.g., using cases to predict hospitalizations) consistently improves Mantis’s accuracy across all forecast horizons. Relative MAE shown for COVID-19 mortality forecasts with (blue) and without (orange) hospitalization covariates across 2, 4, 6, and 8-week horizons. (b) Mantis’s prediction interval widths increase appropriately with forecast horizon, reflecting growing uncertainty in longer-range predictions. The 50% confidence interval width (blue) grows from approximately 50 to 80 deaths, while the 95% interval width (orange) expands from approximately 110 to 180 deaths over 1 to 4 weeks ahead when forecasting COVID-19 mortality. This systematic widening demonstrates the model becoming appropriately less confident as prediction horizon increases.
@@
We evaluated Mantis’s long-range forecasting performance at an 8-week horizon across the same benchmark tasks and baselines described above. Mantis consistently achieved the lowest relative MAE across all diseases and targets tested, demonstrating its ability to extrapolate epidemiological dynamics further into the future than prior models. At the 8-week horizon, Mantis achieved lower relative error than all models in the CDC Forecast Hub at 4 weeks, and lower raw error than the median 4-week error of Forecast Hub models, demonstrating long-range accuracy that exceeds the short-term limits of existing expert-tuned systems.
@@
Figure 3: Mantis Produces Accurate and Generalizable Forecasts Across Diseases and Geographies. (a) Four-week-ahead forecasts (blue dashed line and shaded 90% CI) compared to observed outcomes (black) for COVID-19 mortality in Minnesota and influenza-like illness (ILI) in Michigan. In the latter, Mantis demonstrates its foundation model capacity by accurately forecasting syndromic inputs despite never being trained with syndromic data. (b) Eight-week-ahead forecasts for four historical outbreaks, highlighting Mantis’s ability to generalize across time, space, and transmission mode. In Alagoas, Brazil, Mantis forecasts the 2004 dengue surge prematurely, highlighting a failure in temporal calibration due to the absence of covariates, but still estimates the peak height with remarkable accuracy despite being 18 weeks early. Forecasts for hepatitis B in Florida—a chronic, bloodborne disease type absent from Mantis’s training set—further demonstrate its capacity to generalize to out-of-distribution disease profiles while maintaining forecast stability and accuracy.
Mantis as a Foundation Model: Generalizing to New Mechanisms and Inputs
@@
First, we benchmarked Mantis on diseases governed by transmission and progression mechanisms never represented in training simulations. For example, hepatitis B exhibits bloodborne transmission and chronic infection dynamics that are markedly different from the acute, self-limiting outbreaks used in pretraining. The Mantis training set included no models without recovery—all simulated diseases eventually transition out of the infectious state—and the longest average infectious period is just 10 days. In contrast, hepatitis B infection can persist for years. Despite this fundamental mechanistic mismatch, Mantis achieved competitive performance on hepatitis B forecasts, ranking second in MAE (0.85 vs 0.82) and achieving the best WIS (0.84) among all models (Table 1). These results demonstrate that Mantis can generalize across fundamental mechanistic boundaries—not just to new pathogens, but to qualitatively different epidemic regimes, such as chronic persistence versus outbreak spread.
Next, we tested Mantis’s ability to generalize across data modalities. Although trained using cases, hospitalizations, and deaths, Mantis can incorporate new input types at inference time. We evaluated Mantis on syndromic surveillance data (percent of outpatient visits for influenza-like illness), a data type never represented in training simulations. Despite this distribution shift, Mantis achieved the lowest MAE (0.83) and WIS (0.88) on ILI forecasting (Table 1), demonstrating robust generalization to novel surveillance indicators. This capability is especially valuable in settings where conventional case counts may be missing or delayed, but syndromic indicators are available in near real-time.

### v4

# Mantis: A Foundation Model for Mechanistic Disease Forecasting

@@ Abstract
Infectious disease forecasting in novel outbreaks or low-resource settings is hampered by the need for large disease and covariate data sets, bespoke training, and expert tuning, all of which can hinder rapid generation of forecasts for new settings. To help address these challenges, we developed Mantis, a foundation model trained entirely on mechanistic simulations, which enables out-of-the-box forecasting across diseases, regions, and outcomes, even in settings with limited historical data. We evaluated Mantis against 48 forecasting models across six diseases with diverse modes of transmission, assessing both point forecast accuracy (mean absolute error) and probabilistic performance (weighted interval score and coverage). Despite using no real-world data during training, Mantis achieved lower mean absolute error than all models in the CDC’s COVID-19 Forecast Hub when backtested on early pandemic forecasts which it had not previously seen. Across all other diseases tested, Mantis consistently ranked in the top two models across evaluation metrics. Mantis further generalized to diseases with transmission mechanisms not represented in its training data, demonstrating that it can capture fundamental contagion dynamics rather than memorizing disease-specific patterns. These capabilities illustrate that purely simulation-based foundation models such as Mantis can provide a practical foundation for disease forecasting: general-purpose, accurate, and deployable where traditional models struggle.
@@ Introduction
Accurate forecasting of infectious disease dynamics is fundamental to effective public health response. During the COVID-19 pandemic, forecasts shaped many aspects of daily life, guiding decisions about how to allocate scarce hospital beds, whether restaurants could reopen, and if children could safely attend school [1]. Despite their importance, however, disease forecasts often struggle when they are most needed. In the first two years of the COVID-19 pandemic, the performance of many models in the CDC Forecast Hub was worse than a naïve baseline that simply projected the most recent value forward [2]. This highlights a key limitation of many operational systems: accurate, actionable forecasts are most difficult when uncertainty is highest, but this is often also when decisions are most urgent.
@@
Several structural limitations contribute to this challenge. First, many models require large volumes of disease-specific historical data, which are rarely available early in outbreaks. Second, they rely on extensive manual tuning by expert teams, limiting scalability and access in under-resourced health departments, such as those in rural counties or low- and middle-income countries. Third, they generalize poorly across diseases, locations, or targets, necessitating new models for every context.
@@
In addition, forecasting horizons are typically relatively short: most models predict 1–4 weeks ahead [2, 3, 4], in some cases too short-term for public health decisions. Key decisions, such as ramping up hospital capacity, acquiring ventilators, or coordinating vaccination campaigns, often require 6–8 weeks of lead time. Without reliable long-range forecasts, these decisions are either delayed or made reactively, often at considerable cost [5]. Typically, projections beyond 4 weeks are restricted to scenario analysis, which explores hypothetical outcomes conditional on specific policy interventions or behavioral changes [6, 7]. However, unconditional forecasts at longer horizons remain valuable for public health planning. We define forecasts as unconditional projections of expected outcomes, versus scenario analyses that specify hypothetical interventions (i.e. conditional projections). By this definition, temporal horizon does not determine whether a projection is a forecast or a scenario—an 8-week unconditional projection is a forecast, while a 2-week conditional projection is still scenario analysis (given that policy changes, new variants, and other scenarios can, in principle, happen on any timeframe).
Addressing the challenges of limited historical data, poor cross-disease generalization, restricted forecast horizons, and the need for extensive expert tuning requires a different approach to model training. Foundation models offer a promising paradigm: by training large neural networks on massive, diverse datasets, these models learn generalizable representations that enable strong out-of-the-box performance across a wide range of tasks without task-specific retraining [8, 9]. This approach has been highly successful across multiple domains, such as natural language processing and computer vision, where a single pretrained model can be deployed immediately for diverse applications [8, 10]. A foundation model for disease forecasting could provide similar benefits: accurate predictions across pathogens, geographies, and targets without the expert tuning and historical data typically required for each new setting. However, the data requirements for foundation models present a fundamental obstacle in epidemiology. Major outbreaks are rare, surveillance systems vary widely across regions and time periods, and no historical dataset can span the full diversity of potential pathogens. Real-world disease data alone cannot provide the scale and coverage needed to train a truly general-purpose model.
Mechanistic simulation offers an alternate approach. Epidemiological models encoding disease transmission, progression, and intervention dynamics can generate unlimited synthetic outbreak data spanning scenarios never observed in practice. By training neural networks on these simulations rather than historical observations, we can achieve the data scale required for foundation models while embedding mechanistic scientific knowledge about contagion dynamics directly into the learning process. We previously developed this approach, called Simulation-Grounded Neural Networks (SGNNs), and demonstrated its effectiveness across multiple scientific domains, including initial tests in epidemiological forecasting [11].
@@
In this study, we present Mantis, a foundation model for disease forecasting using the SGNN framework. We expand our training simulations to synthetic outbreaks from mechanistic models that capture transmission, progression, immunity, seasonality, and intervention effects. We simulate human-to-human disease spread, as well as environmental and vector-borne diseases. These simulations encode core epidemiological mechanisms, allowing Mantis to internalize general contagion dynamics rather than purely extrapolating from historical trends. Unlike traditional models, Mantis requires no disease-specific tuning or retraining, but still produces accurate probabilistic forecasts out-of-the-box across diseases, geographies, and targets (Figure 1).
@@
Figure 1: Conceptual overview of Mantis. Mantis is a simulation-grounded foundation model trained entirely on synthetic outbreaks generated by mechanistic epidemiological models. The training pipeline begins with a modular simulator that encodes diverse outbreak mechanisms, including multiple transmission modes (human-to-human, vectorborne, environmental), progression dynamics, intervention strategies, and population structures. These simulations are then passed through an observation model that incorporates real-world surveillance effects such as underreporting, stochasticity, and reporting delays. Trained on over 400 million simulated days, Mantis forecasts directly from real-world time series at inference time. Mantis enables accurate out-of-the-box performance and longer-range forecasting capability.
@@
We evaluated Mantis against 48 models across forecasting tasks spanning six real-world diseases: seasonal influenza, early COVID-19, dengue, hepatitis B, smallpox, and scarlet fever. These evaluations cover multiple transmission modes, outbreak phases, surveillance systems, and historical eras. We assessed both point forecast accuracy and uncertainty quantification. When benchmarked against CDC Forecast Hub models on early pandemic COVID-19 forecasts, Mantis achieved the lowest mean absolute error despite using no COVID-specific data or covariates in training and was tied for the second lowest weighted interval score. Across all other diseases—including those with transmission mechanisms absent from its training data—Mantis consistently ranked among the top two models on all evaluation metrics, demonstrating robust generalization to diverse epidemiological settings.
@@
To support deployment, we release Mantis as an open-source repository. Users can generate forecasts from a single historical time series without any machine learning expertise.222github.com/carsondudley1/Mantis: Pretrained models, example data, and an interactive forecasting demo are available. Mantis is currently contributing forecasts to a new local-scale forecasting hub within Michigan for the 2025-2026 respiratory virus season and will be submitted to the CDC Forecast Hub to support real-time public health decision-making.
@@ Results
We compared Mantis against a range of published and re-implemented models, including top submissions to the CDC COVID-19 Forecast Hub, statistical baselines (exponential smoothing and SARIMA), LSTM neural networks retrained at each forecast point, and Chronos, a state-of-the-art time-series foundation model [20].
Mantis achieved competitive performance across diseases and metrics (Table 1). For early pandemic COVID-19 forecasts at the state level, Mantis achieved the lowest MAE among evaluated models, including the CDC Forecast Hub ensemble that combines dozens of expert-tuned submissions. The CDC ensemble achieved slightly better WIS (0.61 vs 0.63), while UMass-MechBayes achieved marginally better 95% coverage (0.93 vs 0.91), demonstrating that multiple approaches can excel on different aspects of probabilistic forecasting. Although results have shown that some general time series forecasting foundation models can confer reasonable results for infectious disease forecasting [21], we did not find that any we tested (Chronos [20] and TimePFN [22]) gave competitive performance with Mantis or our other baseline models, so we did not include them in our analysis.
@@
Example forecast trajectories illustrate Mantis’s ability to match the timing and magnitude of real-world outbreaks across epidemiological settings (Figure 3, with over 100 more in the supplement). Forecasts include uncertainty intervals derived from the model’s predicted quantiles, enabling direct use in probabilistic planning. Because Mantis is trained entirely on broad mechanistic simulations rather than historical observations, it requires no model fitting or fine-tuning for new targets.
@@
State-level COVID-19 [14] Mantis 0.65 0.63 0.51 0.91
@@
Local-scale flu hosp. [12] Mantis 0.85 0.86 0.51 0.92
@@
Dengue cases [15] Mantis 0.75 0.80 0.47 0.90
@@
Smallpox cases [17] Mantis 0.83 0.84 0.45 0.87
@@
Scarlet Fever cases [18] Mantis 0.80 0.81 0.47 0.91
@@
Hepatitis B cases [16] Mantis 0.85 0.84 0.47 0.90
@@
We evaluated two variants of Mantis to assess the impact of additional information on forecasting performance: a covariate-free version, which receives only the time series to be forecast (e.g., deaths), and a covariate-aware version, which also receives related signals (e.g., cases or hospitalizations). (See Methods for training details for each model.)
For example, when forecasting COVID-19 deaths, the covariate-aware Mantis version (used in evaluations) receives both historical deaths and upstream hospitalization data, while the covariate-free version sees only the death time series. Both models use identical architecture and training data; the only difference is the presence or absence of additional input channels.
@@
Figure 2: Mantis performance and uncertainty across forecast horizons with and without covariates. (a) Covariate integration improves accuracy. Including covariates (e.g., using cases to predict hospitalizations) consistently improves Mantis’s accuracy across all forecast horizons. Relative MAE vs. naive baseline shown for COVID-19 mortality forecasts with (blue) and without (orange) hospitalization covariates across 2, 4, 6, and 8-week horizons. (b) Mantis maintains calibrated uncertainty across forecast horizons. Mantis’s prediction interval widths increase appropriately with forecast horizon, reflecting growing uncertainty in longer-range predictions. The 50% confidence interval width (blue) grows from approximately 50 to 80 deaths, while the 95% interval width (orange) expands from approximately 110 to 180 deaths over 1 to 4 weeks ahead when forecasting COVID-19 mortality. This systematic widening demonstrates the model becoming appropriately less confident as prediction horizon increases.
@@
We evaluated Mantis’s long-range forecasting performance at an 8-week horizon across the same benchmark tasks and baselines described above. Mantis consistently achieved the lowest relative MAE across all diseases and targets tested, demonstrating its ability to extrapolate epidemiological dynamics further into the future than prior models. At the 8-week horizon, Mantis achieved lower relative error than all models in the CDC Forecast Hub at 4 weeks, and lower raw error than the median 4-week error of Forecast Hub models, demonstrating long-range accuracy that exceeds the short-term limits of some existing systems.
@@
Figure 3: Mantis Produces Accurate and Generalizable Forecasts Across Diseases and Geographies. (a) Four-week-ahead forecasts (blue dashed line and shaded 90% CI) compared to observed outcomes (black) for COVID-19 mortality in Minnesota and influenza-like illness (ILI) in Michigan. In the latter, Mantis demonstrates its foundation model capacity by accurately forecasting syndromic inputs despite never being trained with syndromic data. (b) Eight-week-ahead forecasts for four historical outbreaks, highlighting Mantis’s ability to generalize across time, space, and transmission mode. In Alagoas, Brazil, Mantis forecasts the 2004 dengue surge prematurely, highlighting a failure in temporal calibration due to the absence of covariates, but still estimates the peak height with high accuracy despite being 18 weeks early. Forecasts for hepatitis B in Florida—a chronic, bloodborne disease type absent from Mantis’s training set—also demonstrate its capacity to generalize to out-of-distribution disease profiles while maintaining forecast stability and accuracy.
Testing Mantis as a Foundation Model: Generalizing to New Mechanisms and Inputs
@@
First, we benchmarked Mantis on diseases governed by transmission and progression mechanisms never represented in training simulations. For example, hepatitis B exhibits bloodborne transmission and chronic infection dynamics that are markedly different from the acute, self-limiting outbreaks used in pretraining. The Mantis training set included no models without recovery—all simulated diseases eventually transition out of the infectious state—and the longest average infectious period is just 10 days. In contrast, hepatitis B infection can persist for years. Despite this fundamental mechanistic mismatch, Mantis achieved competitive performance on hepatitis B forecasts, ranking second in MAE (0.85 vs 0.82) and achieving the best WIS (0.84) among models we tested (Table 1). These results demonstrate that Mantis can generalize across mechanistic boundaries—not just to new pathogens, but to qualitatively different epidemic regimes, such as chronic persistence versus outbreak spread. However, it is important to note that this forecasting task is limited: hepatitis B data exhibit relatively stable but noisy dynamics without substantial outbreak structure. While Mantis performs well compared to baselines, there is relatively little signal to exploit beyond tracking a noisy equilibrium, and 8-week forecasts of endemic incidence are rarely a high-priority public health need. The point of this benchmark is not to demonstrate operational utility for hepatitis B, but to show that Mantis generalizes to qualitatively different epidemic regimes rather than memorizing a narrow set of outbreak patterns from training. However, given the relatively simple dynamics here, further testing of this point would be a useful step for future work.
Next, we tested Mantis’s ability to generalize across data modalities. Although trained using cases, hospitalizations, and deaths, Mantis can incorporate new input types at inference time. We evaluated Mantis on syndromic surveillance data (percent of outpatient visits for influenza-like illness), a data type never represented in training simulations. Despite this distribution shift, Mantis achieved the lowest MAE (0.83) and WIS (0.88) on ILI forecasting among models we tested (Table 1), demonstrating generalization to novel surveillance indicators. This capability is especially valuable in settings where conventional case counts may be missing or delayed, but syndromic indicators are available in near real-time.
@@
Ablation Studies: The Role of Scale and Mechanistic Pretraining
To understand the key factors driving Mantis’s performance, we conducted two ablation studies examining the roles of architectural capacity and mechanistic pretraining in enabling accurate out-of-the-box forecasting.
Architectural Capacity Requirements
Despite identical training data and procedures, the reduced architectural capacity led to substantially degraded performance. When evaluated on COVID-19 mortality forecasting across U.S. states, the LSTM architecture achieved mean absolute error (MAE) that was 89% higher than the full Mantis model. This demonstrates that while synthetic training is useful, it requires sufficient architectural capacity to effectively learn and generalize from the diverse mechanistic patterns encoded in the simulation data.
Mechanistic Pretraining is Essential for Novel Outbreak Performance
To isolate the contribution of mechanistic pretraining, we trained the full Mantis architecture exclusively on real-world data, eliminating simulations from the training process. We evaluated on COVID-19 death forecasting. We implemented a rolling training strategy that provided the strongest possible baseline: the model was initially trained on the first few months of pandemic data, then progressively retrained with expanding datasets as more real-world observations became available. This approach mimics operational practice and ensures that the real-world-only model has access to all available historical information at each forecast point.
The results reveal the critical importance of mechanistic pretraining. When trained only on real-world COVID-19 data, forecast error increased by 262% compared to the simulation-pretrained Mantis model. This substantial performance degradation reflects the fundamental challenge of learning outbreak dynamics from limited historical data during a novel pandemic. Real-world training data, even when accumulated over months, cannot capture the full spectrum of epidemic behaviors that might emerge. In contrast, the simulation-pretrained model draws from hundreds of millions of days of diverse outbreak scenarios, enabling it to recognize and extrapolate patterns even in unprecedented contexts.
These ablation results underscore the core insight behind Mantis: architectural capacity provides the necessary capability to represent complex patterns, while large-scale simulation-grounded training paradigm provides abundant data to learn generalizable principles of epidemic dynamics and utilize the enhanced architectural capacity. Neither component is sufficient in isolation. Rather, the combination of high-capacity modern neural architectures and large-scale mechanistic simulations enables out-of-the-box generalization to novel diseases and outbreak contexts.